OpenStudy (anonymous):

csc^2-cosx secx=cot^2

OpenStudy (xapproachesinfinity):

first you have to use this 1+cot^2x=csc^2x

OpenStudy (xapproachesinfinity):

then observe what you can do next

OpenStudy (anonymous):

Or, starting from the left side, \[\cos x\sec x=\frac{\cos x}{\cos x}=1\] so you have \[\csc^2x-1=\cot^2x\]